201210 R1ProgressRep

From New IAC Wiki
Jump to navigation Jump to search


Sector 1 - began training stringers to turn on high voltage and look for signals using STB probe and ADB monitoring system.

Sector 5 - restrung Guard and Field wires for replaced feed throughs in SL1 (Fields: 167-11, 187-2, 223-8 and 225-11. Guards: 227-10 and 228-3). All wires strung were continuity tested for potential crossed wires.

Sector 6 - tested SL2 Field wires from row 134 to 168 (245 wires).


Sector 1 - Injected a 10 MHz square wave with a 40 mV peak-peak amplitude into each channel of the ADB monitor system. Tested the gain on A and B of each channel. Preliminary results show 5 channels that may not be functioning properly.

Sector 5 - removed and replaced 4 more damaged or glue contaminated feed throughs that were found during visual inspection (SL1: 16-2 (Sense). SL2: 37-1 (Guard), 227-8 (Field) and 227-9 (Field)).

Sector 6 - tested SL2 Field wires from row 169 to 189 (147 wires).

Temperatures in the cleanroom have been decreasing slowly over the last week. Increased the temperature set point of the HVAC system by one degree to keep temperatures within cleanroom specifications. Will monitor and report any issues.


Sector 1 - Injected a 1kHz -40mV peak pulse with a 0.1% duty cycle directly into the ADB monitor system to verify ECL output for all 96 channels. All channels confirmed a -1V peak with a 1us pulse width. Mapped Layer/Wire to corresponding ADB channel. The map that was sent to us does not correspond between ADB channel and Layer/Wire.

Sector 6 - tested Field wires from row 190 to 227 (263 wires).


Sector 1 - Spent considerable time trouble shooting ribbon cables to verify that an injected signal maps correctly from each STB channel to each ADB channel and produces an ECL output.

Sector 5 - completed weekly Tension Trend test.

Sector 6 - completed weekly Tension Trend test. Tested SL2 Sense wires from row 3 to 39 (109 wires).

Completed weekly clean in cleanroom.


Sector 1 - reorganized rack to accommodate for the VME crate.

Sector 5 - rotated chamber; repositioned magnet; leveled the stringing arm and aligned the chamber for Sense wire stringing in SL2.

Sector 6 - tested SL2 Sense wires from row 40-146 (321 wires).


Sector 1 - continuity tested all STB to ADB cables. R1SC4 and R1SC6 appear to be the only completely functioning cables.

Sector 5 - finished removing debris from SL2.

Sector 6 - rotated chamber; repositioned magnet and leveled stringing arm. Will align the chamber for SL2 stringing tomorrow.


Sector 1 - DC tested STB to ADB cables to check to see if replacing end connectors would fix the 4 bad cables.

Sector 5 - continuity tested all remaining Field wires. Removed a few more pieces of metal debris that were found inside the chamber.

Sector 6 - aligned the chamber for SL2 wire replacement for feed throughs that were removed and replaced. Replaced and tested wires and began final visual inspection before potting begins next week.


Sector 1 - worked on replacing connector ends on STB to ADB cable.

Sector 5 - restrung Field and Guard wires from replaced feed throughs (SL2: 227-8F, 227-9F, 37-1G, 228-3G). Completed weekly Tension Trend test.

Sector 6 - completed weekly Tension Trend test. Confirmed tension values for 3 wires that show inconsistent behavior in the Tension Trend (SL1: 90-2S. SL2: 20-7F, 60-7F). All three have a crimp pin that was easily moved in the feed through. Crimp pins were seated well and final tension measurements were taken. All are within tension specifications. (The inconsistent behavior can be easily visualized with the DBplotter.)

Completed weekly cleanroom clean.


Sector 1 - Injected a 40 KHz -40 mV peak amplitude pulse into HVTB, connected HVTB to ADB with cables that passed pin-to-pin wire continuity checks with no cross-continuity with other channels (R1 SC4, R1 SC6) and tested ECL output of the ADB from the injected signal. -Found cable-strain at the connection point on the HVTB was the cause of excessive noise we observed from the amplified ADB signal that was probed from the front panel ports. Slight adjustments of the signal cable stabilize the signal, but if the signal cables ever get bumped the signal goes bad easily and the cables have to be re-adjusted. This problem may be an indication the cable needs to be replaced, or that we need to make sure there's minimal cable strain at the HVTB port. - With the first probe we used (a 2-pin connector, with one pin grounded) ECL output had significant noise, but the logic signal was present. Switching to a passive 1x grounded scope probe eliminated the noise. We observed ECL pulses at an interval of ~1 μs, width 200ns, peak ~150 mV that aren't triggered by the injected signal. Logic pulses that were in parallel (seem to be triggered by injected pulse) had a wider ~1 μs pulse and 150 mV peak. Switching pins on a given channel does reverse the polarity of the peak (positive vs. negative) as one would expect. We have only tested channels 1 and 2 of the ADB and HVTB so far.

Sector 5 - began stringing Sense wires in SL2. Strung 40 Sense wires (now on SL2 - row 17).

Sector 6 - Marcie is working on compiling a QC document to record any and all issues, actions, etc. that occurred to the chamber during stringing. Pre-potting tension testing is complete. Chamber data is being reviewed and potting will commence once all concerned parties have signed off on the tension data.


Just after noon today, the HVAC system went down. All stringing activities were immediately halted and wire, crimp pins, etc. were covered. Upon investigation, it was determined that the blower motor belt had broken. A work order was placed and ISU Facilities replaced the belt at approximately 3:45pm. Temperature and humidity in the cleanroom increased to 75° F and 40%, respectively. Temperatures and particles will be monitored until levels are within specifications before any stringing activities resume.

There was a power bump at 1:25pm that resulted in spikes on the temperature and humidity monitoring graphs. Temperature and humidity levels appear to be recording normally now, although the graphs look distorted. The graphs should appear normal after the spike readings have moved beyond the monitoring periods of each chart. (Temperature: http://hal.iac.isu.edu/cacti-graphs/graph_6.html ) (Humidity: http://hal.iac.isu.edu/cacti-graphs/graph_7.html ).

At approximately 4:15pm cleanroom temperatures dropped below the upper limit of 72° F. The humidity level at that time was 36%.

Sector 1 - began spot testing ECL output on ADB. Changed op-amp on channel 90. Attempted trouble shooting to eliminate noise throughout the system.

Sector 5 - strung 33 Sense wires (now on SL2 - row 28).

Sector 6 - another Field glue contaminated feed through was found (SL2: 28-4 top).

At 5:52pm the last recorded temperature readings by probe number are as follows: (1)71.20°, (2)69.90°, (3)70.20° and (4)69.90°. The humidity level was at 41.20%.


Temperatures seem to have stabilized. At 16:45, the daily 5 minute average for each probe was as follows: (1)70.60°, (2)70.17°, (3)70.18°, (4)70.17°. The humidity daily 5 minute average for the same time was 28.66%.

Sector 1 - still trouble shooting to identify noise sources within the system. Started taking screen shots of signals from ECL and ADB front ports.

Sector 5 - strung 54 Sense wires (now on SL2 - row 46).


Temperatures are still within specifications. At 17:45, the daily 5 minute average for each probe was as follows: (1)70.49°, (2)70.14°, (3)70.12°, (4)70.11°. The humidity daily 5 minute average for the same time was 21.05%.

Sector 1 - Identified cross-talk occurring on wires/channels near the channel that is having a signal injected. Obtained a new signal cable for HVTB and ADB to try on Monday. Signal cables sent to us use connectors that do not provide strain relief for the knife contacts with the wires in the cable. Had one pull apart, also identified this as a source of poor contact points within the wire that was causing significant amounts of ringing in the ADB-front panel's analog output signals. New cable should take care of this problem. Analyzed output of Ch. 1-14, characterized behavior with existing bad cabling. Also put the "bad" op-amp we identified from Ch. 90 on the HVTB into Ch. 96 and observed ADB and ECL signals with a "good" and "bad" op-amp. Currently the good op-amps in all channels result in two "ragged" pulses per injected pulse in the ADB-front panel's analog signals and variable triggering behavior with the ECL that would result in junk data if hooked up to a TDC. The "ragged" pulses are most likely an artifact of the cabling being used, and should be resolved with the new cabling.

Sector 5 - strung 45 Sense wires (now on SL2 - row 61). Completed weekly Tension Trend test.

Sector 6 - completed weekly Tension Trend test.

Conducted a weekly clean of the cleanroom.


Sector 5 - strung 24 Sense wires (now on SL2 - row 61).


Sector 1 - Changes in experimental setup:

-Trying a new 34-pin ribbon cable for getting signal from HVTB to ADB

-Using a second probe for the front-panel ADB Analog output ports (Ch. 1 Yellow)

(Orientation of the ribbon and/or the particular channel being probed on the ADB seem to have a distinct effect on the quality of the signal observed. Two configurations were tried today.)

Sector 5 - strung 41 Sense wires (now on SL2 - row 83).

Sector 6 - pulled and replaced two more feed throughs (SL1: 130-4F top and SL2: 28-4F top).


Sector 5 - strung 51 Sense wires (now on SL2 - row 100).

Sector 6 - restrung wires in replaced feed throughs in both superlayers. Tested those wires and all adjacent wires. Completed continuity testing. Updated pre-potting tension spreadsheets.


Sector 5 - strung 42 Sense wires (now on SL2 - row 114). Completed weekly Tension Trend Test.

Sector 6 - completed weekly Tension Trend Test.

Completed weekly cleanroom clean.


Conducted potting training with all stringers.

Sector 5 - strung 21 Sense wires (now on SL2 - row 121). Had issues with the left roller on the stringing machine. Replaced balloon but the roller is still not pulling wire through correctly. Tried using one of the rollers from the Sector 6 stringing machine. Still having issues with the wire feed. Will trouble-shoot tomorrow.

Sector 6 - removed stringing rail, etc. in preparation for potting. Began potting top plate in SL2. Completed from row 1 to row 62.


Sector 5 - strung 51 Sense wires (now on SL2 - row 138). Stringing machine issues were fixed by replacing rollers with those from the Sector 6 stringing machine. If the Sector 5 rollers can not be repaired, we may need to explore options for replacement. The soft, wool like material that surrounds the rollers is not uniform and causes an uneven surface once the roller is covered with the balloon.

Sector 6 - potted top plate in SL2 from row 63 - 141.


Sector 5 - strung 63 Sense wires (now on SL2 - row 159).

Sector 6 - completed potting top plate in SL2. Started potting top plate in SL1 from row 228 - 135.